# Counting and Entropy Bounds for Structure-Avoiding Spatially-Coupled LDPC Constructions

### v1

@@ Abstract
Designing large coupling memory quasi-cyclic spatially-coupled LDPC (QC-SC-LDPC) codes with low error floors requires eliminating specific harmful substructures (e.g., short cycles) induced by edge spreading and lifting. Building on our work [8] that introduced a Clique Lovász Local Lemma (CLLL)-based design principle and a Moser–Tardos (MT)-type constructive approach, this work quantifies the size and structure of the feasible design space. Using the quantitative CLLL, we derive explicit lower bounds on the number of partition matrices satisfying a given family of structure-avoidance constraints, and further obtain bounds on the number of non-equivalent solutions under row/column permutations. Moreover, via Rényi-entropy bounds for the MT distribution, we provide a computable lower bound on the number of distinct solutions that the MT algorithm can output, giving a concrete diversity guarantee for randomized constructions. Specializations for eliminating 4-cycle candidates yield closed-form bounds as functions of system parameters, offering a principled way to size memory/lifting and to estimate the remaining search space.
@@ I Introduction
Algebraic constructions can deterministically control girth and certain substructures, but they are typically restricted to narrow templates and thus do not readily yield explicit lower bounds on how many SC-LDPC code realizations satisfy prescribed structure-avoidance constraints. In contrast, randomized/heuristic design explores a much larger space—e.g., the OO–CPO framework [2] (extended in [6]) and the probabilistic optimization approach [18]—yet its search runtime can be highly instance-dependent; our Moser–Tardos (MT) resampling approach in [8] can be viewed as a randomized OO–CPO special case that locally resamples partitioning/lifting variables whenever an OO/CPO-type constraint is violated.
In this paper, we extend our existence and constructive results in [8] from feasibility to quantitative guarantees on the solution-space size, by deriving explicit counting and entropy bounds for structure-avoiding SC-LDPC code constructions.
Contributions: (i) Under joint non-uniform edge spreading and non-uniform lifting, we apply a quantitative clique LLL to derive explicit lower bounds on the number of structure-avoiding QC-SC-LDPC designs, including bounds on non-equivalent solutions under row/column permutations; (ii) we lower bound the number of distinct solutions producible by the MT algorithm via Rényi-entropy bounds of the MT output distribution, with closed-form specializations for eliminating $4$-cycle candidates; (iii) we connect the MT resampling construction to the OO–CPO design framework by proving that circulant power optimization (CPO) subsumes MT as a special case (Proposition 2); consequently, the output-diversity guarantee in (ii) provides theoretical evidence for the strong potential of CPO-based QC-SC-LDPC design procedures [2, 6, 18].
@@ II-B Major Harmful Structures
This naturally motivates the common two-stage design strategy: first choose $\mathbf{P}$ to suppress protograph cycle candidates, then select $\mathbf{L}$ to further eliminate cycles (and induced harmful substructures) after lifting, as in [2, 18]. In this work we focus on the AWGN setting, where dominant error-floor mechanisms are often attributed to low-weight absorbing sets (ASs) [2].
@@ II-C Constraint Satisfaction Problem
$X_{i,j}$ $=(P(i,j),L(i,j))\in\{0,\ldots,m\}\times\mathbb{Z}_{Z}.$
For each $H_{\ell}\in\mathcal{H}$, define a constraint (bad event) $A_{\ell}$ that $H_{\ell}$ becomes active after edge spreading and $Z$-lifting. Then finding a feasible pair $(P,L)$ that avoids all $H_{\ell}$ is exactly the CSP
@@
A CSP is specified by (i) a set of variables and (ii) a set of constraints on these variables. Here the variables are the $\gamma\kappa$ edge-variables $X_{i,j}=(P(i,j),L(i,j))$. Each harmful structure $H_{\ell}$ induces a local “activation” condition after edge spreading and lifting, so we impose the constraint $\overline{A_{\ell}}$ that forbids this activation. Importantly, whether $A_{\ell}$ occurs depends only on the entries of $(P,L)$ on the edges participating in $H_{\ell}$ (for cycle candidates this is exactly the local conditions (1) and (2) in Lemma 1). Hence each constraint $\overline{A_{\ell}}$ involves only a subset of variables, and the feasibility problem $\bigcap_{\ell=1}^{n}\overline{A_{\ell}}$ is precisely a CSP. The bipartite graph $G_{E}$ is then formed by linking $A_{\ell}$ to the variables it depends on, as in Sec. II-C. ∎
Building on our earlier work [8], we continue to formulate structure-avoiding QC-SC-LDPC code design as a CSP and employ CLLL–based tools to characterize ensembles relevant to the high-SNR (error-floor) regime. In contrast to [8], we further quantify the feasible design space via counting and entropy bounds.
@@ III-A Lower Bound on the Number of Solutions
Under the CSP formulation in Proposition 1, each QC-SC-LDPC design corresponds to an assignment of the $\gamma\kappa$ edge variables $X_{i,j}=(P(i,j),L(i,j))$, i.e., a pair $(P,L)$ consisting of an edge-spreading (partition) matrix $P$ and a lifting-shift matrix $L$. Therefore, using the quantitative version of the LLL, we can not only obtain existence results but also establish an explicit lower bound on the number of feasible pairs $(P,L)$ that satisfy all structure-avoidance constraints, thereby quantifying the size of the feasible solution space. The size and structural information of this solution space are crucial for finding codes with good overall performance at lower complexity. In this subsection, we first use the quantitative LLL to derive a theoretical lower bound on the number of feasible solutions $(P,L)$, and then use an entropy method (via Rényi-entropy bounds for the MT output distribution) to obtain a computable lower bound on the number of feasible pairs $(P,L)$ that the algorithm can output.
@@
For any QC-SC-LDPC code constructed with respect to a $\gamma\times\kappa$ fully-connected block code base graph, with no parallel edges, under the coupling pattern $\mathbf{a}=(a_{0},a_{1},\ldots,a_{m_{t}})$ and corresponding (non-uniform) distribution $\mathbf{p}=(p_{0},p_{1},\ldots,p_{m_{t}})$, and under the lifting-shift alphabet $\mathbf{b}=(b_{0},b_{1},\ldots,b_{Z-1})=(0,1,\ldots,Z-1)$ with distribution $\mathbf{q}=(q_{0},q_{1},\ldots,q_{Z-1})$.Assume the random construction chooses each $P(i,j)$ independently with $\Pr[P(i,j)=\ell]=p_{\ell}$, and chooses each lifting shift $L(i,j)$ independently with $\Pr[L(i,j)=b_{t}]=q_{t}$, and $P$ is independent of $L$.
@@
Here, $(\bigcap_{i\in[k]}\bar{H_{i}})$ denotes the number of feasible assignments of $(P,L)$ (equivalently, solutions in the CSP variable set $\chi$) that make $\bigcap_{i\in[k]}\bar{H_{i}}$ true.
@@
Let $\Omega:=\big(\{0,1,\ldots,m_{t}\}\times\mathbb{Z}_{Z}\big)^{\gamma\kappa}$ be the assignment space of all edge-variables $X_{i,j}=(P(i,j),L(i,j))$, and let
$S:=\bigcap_{i\in[k]}\bar{H_{i}}$
$\Pr_{\Omega}(\omega)$ $=\prod_{i\in[\gamma],\,j\in[\kappa]}\Pr[P(i,j)=\omega^{P}_{i,j}]\Pr[L(i,j)=\omega^{L}_{i,j}]$
$\leq(p_{\max}q_{\max})^{\gamma\kappa}.$
Hence,
$\Pr_{\Omega}(S)=\sum_{\omega\in S}\Pr_{\Omega}(\omega)\leq\#(S)\cdot(p_{\max}q_{\max})^{\gamma\kappa},$
which implies
$(S)\geq\left(\frac{1}{p_{\max}q_{\max}}\right)^{\gamma\kappa}\Pr_{\Omega}(S).$ (4)
By the quantitative clique-LLL bound used in [8, Lemma 2] (under the stated condition on $\max_{i}P_{H_{i}}$),
$\Pr_{\Omega}(S)\geq\prod_{v\in[m]}\Big(1-\sum_{i\in K_{v}}x_{i,v}\Big).$
Combining with (4) yields
$(S)\geq\left(\frac{1}{p_{\max}q_{\max}}\right)^{\gamma\kappa}\prod_{v\in[m]}\Big(1-\sum_{i\in K_{v}}x_{i,v}\Big).$
Finally, substituting into [8, Theorem 1] the same choice of $\{x_{i,v}\}$ that yields the two explicit cases $I$ and $II$ gives the claimed bound.
@@
${\#\Big(\bigcap_{c_{4}\in\mathcal{C}_{4}}\bar{c_{4}}\Big)}\geq[Z(m_{t}+1)]^{\gamma\kappa}\begin{cases}(1-\frac{2}{\Delta})^{\binom{\gamma}{2}\binom{\kappa}{2}},&{\text{if}}\ I>II,\\[2.0pt] \Big(1-\frac{\gamma\kappa-\gamma-\kappa+1}{4(\gamma\kappa-\gamma-\kappa)}\Big)^{\gamma\kappa},&{\text{otherwise.}}\end{cases}$
@@
We use the notion of equivalence in this paper to highlight that neither decoding threshold nor the number of combinatorial objects, e.g., cycles, absorbing sets [25], trapping sets [26], etc., change with a sequence of row and column permutations of a binary matrix.
Lemma 2 derives the congruence between coupled proto-matrices and the partitioning matrices that are used to construct them. This congruence allows searching for a coupled SC-LDPC code over a reduced search space of small $(\gamma\times\kappa)$ nonequivalent partitioning matrices.
@@
For any QC-SC-LDPC code constructed with respect to a $\gamma\times\kappa$ fully-connected block code base graph, with no parallel edges, under the coupling pattern $\mathbf{a}=(a_{0},a_{1},...,a_{m_{t}})$ and corresponding probability distribution $\mathbf{p}=(p_{0},p_{1},...,p_{m_{t}})$. If $\{H_{1},...,H_{k}\}$ are a series of avoidable harmful structures in the base graph with dependency graph $G$, and $\{P_{H_{1}},...,P_{H_{k}}\}$ are the probabilities that they are active after the random edge-spreading process, if
@@
${\#_{N}(\bigcap_{c_{4}\in\mathcal{C}_{4}}\bar{c_{4}})}\geq\begin{cases}[\frac{Z(m_{t}+1)]^{\gamma\kappa}(1-\frac{2}{\Delta})^{\binom{\gamma}{2}\binom{\kappa}{2}}}{(\gamma!\kappa!)},&{\text{if}}\ I>II\\ {\frac{[Z(m_{t}+1)]^{\gamma\kappa}(1-\frac{\gamma\kappa-\gamma-\kappa+1}{4(\gamma\kappa-\gamma-\kappa)})^{\gamma\kappa}}{(\gamma!\kappa!)},}&{\text{otherwise.}}\end{cases}$
@@ III-B Diversity of Constructive Outputs via MT Entropy
Under the OO-CPO framework, the CPO operation is closely related to the MT algorithm.
@@
In the OO-CPO framework, circulant power optimization (CPO) includes the Moser–Tardos (M–T) algorithm as a special case.
@@
In the sequel, we show that the MT algorithm can generate a provably large set of distinct feasible solutions. This observation suggests that CPO-based design procedures, such as those in [8]–[10], also have significant potential, since Proposition 2 establishes that, within the OO–CPO framework, circulant power optimization (CPO) subsumes the MT algorithm as a special case.
@@
For any QC-SC-LDPC code constructed with respect to a $\gamma\times\kappa$ fully-connected block code base graph,with no parallel edges, under the coupling pattern $\mathbf{a}=(a_{0},a_{1},...,a_{m_{t}})$ and corresponding probability distribution $\mathbf{p}=(p_{0},p_{1},...,p_{m_{t}})$. If $\{H_{1},...,H_{k}\}$ are a series of avoidable harmful structures in the base graph with dependency graph $G$, and $\{P_{H_{1}},...,P_{H_{k}}\}$ are the probabilities that they are active after the random edge-spreading process, if there is a weighting function $\tilde{\mu}:\mathcal{A}\rightarrow[0,\infty)$ satisfies
@@
for any $i\in[n]$, then the MT Algorithm described above resamples an event $A\in\mathcal{A}$ at most an expected $\tilde{\mu}(A_{i})$ times before it finds such an evaluation. The number of different partition matrices that the algorithm can output, denoted as $_{MT}(\bigcap_{i\in[k]}\bar{H_{i}})$, satisfies the following condition:
@@
where $\Delta=(2\gamma-3)(2\kappa-3)$ , $Z$ is the lifting degree. then the number of different partition matrices that the algorithm can output, denoted as $_{MT}(\bigcap_{c_{4}\in\mathcal{C}_{4}}\bar{c_{4}})$, satisfies the following condition:
@@
see Appendix B. ∎
@@
The number $_{MT-N}(\bigcap_{i\in[k]}\bar{H_{i}})$ of non-equivalent partition matrices satisfying Lemma 1 that can be output by the algorithm is given by:
@@
The number $_{MT-N}(\bigcap_{c_{4}\in\mathcal{C}_{4}}\bar{c_{4}})$ of non-equivalent partition matrices satisfying Lemma 1 that can be output by the algorithm is given by:

### v2

@@ Abstract
Designing large coupling memory quasi-cyclic spatially-coupled LDPC (QC-SC-LDPC) codes with low error floors requires eliminating specific harmful substructures (e.g., short cycles) induced by edge spreading and lifting. Building on our work [9] that introduced a Clique Lovász Local Lemma (CLLL)-based design principle and a Moser–Tardos (MT)-type constructive approach, this work quantifies the size and structure of the feasible design space. Using the quantitative CLLL, we derive explicit lower bounds on the number of feasible edge-spreading and lifting assignments satisfying a given family of structure-avoidance constraints, and further obtain bounds on the number of non-equivalent solutions under row/column permutations. Moreover, via Rényi entropy bounds for the MT distribution, we provide a computable lower bound on the number of distinct solutions that the MT algorithm can output, giving a concrete diversity guarantee for randomized constructions. Specializations for eliminating 4-cycles yield closed-form bounds as functions of system parameters, offering a principled way to select the memory and lifting degree and to estimate the remaining search space.
@@ I Introduction
Algebraic constructions can deterministically control girth and certain substructures, but they are typically restricted to narrow templates and thus do not readily yield explicit lower bounds on how many SC-LDPC code realizations satisfy prescribed structure-avoidance constraints. In contrast, randomized/heuristic design explores a much larger space—e.g., the optimal overlap-circulant power optimization (OO-CPO) framework [3] (extended in [7]) and the probabilistic optimization approach [24]—yet its search runtime can be highly instance-dependent; our Moser-Tardos (MT) resampling approach in [9] can be viewed as a randomized OO-CPO special case that locally resamples partitioning/lifting variables whenever a constraint is violated.
In this paper, we extend our existence and constructive results in [9] from feasibility to quantitative guarantees on the solution-space size, by deriving explicit counting and entropy bounds for structure-avoiding SC-LDPC code constructions. Due to space constraints, full proofs and detailed discussions are deferred to [10].
Contributions: (i) Under joint non-uniform edge spreading and non-uniform lifting, we apply a quantitative clique LLL to derive explicit lower bounds on the number of structure-avoiding QC-SC-LDPC designs, including bounds on non-equivalent solutions under row/column permutations; (ii) we lower bound the number of distinct solutions producible by the MT algorithm via Rényi entropy bounds of the MT output distribution, with closed-form specializations for eliminating $4$-cycles; (iii) we connect the MT resampling construction to the OO-CPO design framework by proving that circulant power optimization (CPO) subsumes MT as a special case (Proposition 2); consequently, the output-diversity guarantee in (ii) provides theoretical evidence for the strong potential of CPO-based QC-SC-LDPC design procedures [3, 7, 24].
@@ II-B Major Harmful Structures
This naturally motivates the common two-stage design strategy: first choose $\mathbf{P}$ to suppress protograph cycle candidates, then select $\mathbf{L}$ to further eliminate cycles (and induced harmful substructures) after lifting, as in [3, 24]. In this work, we focus on the AWGN setting, where dominant error-floor mechanisms are often attributed to low-weight absorbing sets (ASs) [3]. Moreover, all harmful structures considered in this paper are avoidable in the sense that they can be excluded by imposing explicit design constraints within our CSP framework.
@@ II-C Constraint Satisfaction Problem
$X_{i,j}$ $=(\mathbf{P}(i,j),\mathbf{L}(i,j))\in\{0,\ldots,m\}\times\mathbb{Z}_{Z}.$
For each $H_{\ell}\in\mathcal{H}$, define a constraint (bad event) $A_{\ell}$ that $H_{\ell}$ becomes active after edge spreading and $Z$-lifting. Then finding a feasible pair $(\mathbf{P},\mathbf{L})$ that avoids all $H_{\ell}$ is exactly the CSP
@@
A CSP is specified by (i) a set of variables and (ii) a set of constraints on these variables. Here the variables are the $\gamma\kappa$ edge-variables $X_{i,j}=(\mathbf{P}(i,j),\mathbf{L}(i,j))$. Each harmful structure $H_{\ell}$ induces a local “activation” condition after edge spreading and lifting, so we impose the constraint $\overline{A_{\ell}}$ that forbids this activation. Importantly, whether $A_{\ell}$ occurs depends only on the entries of $(\mathbf{P},\mathbf{L})$ on the edges participating in $H_{\ell}$ (for cycle candidates this is exactly the local conditions (1) and (2) in Lemma 1). Hence each constraint $\overline{A_{\ell}}$ involves only a subset of variables, and the feasibility problem $\bigcap_{\ell=1}^{n}\overline{A_{\ell}}$ is precisely a CSP. The bipartite graph $G_{E}$ is then formed by linking $A_{\ell}$ to the variables it depends on, as in Sec. II-C. ∎
Building on our earlier work [9], we continue to formulate structure-avoiding QC-SC-LDPC code design as a CSP and employ CLLL–based tools to characterize ensembles relevant to the high-SNR (error-floor) regime. The method for utilizing CLLL-based tools within the CSP framework is detailed in Appendix F of [10]. In contrast to [9], we further quantify the feasible design space via counting and entropy bounds.
@@ III-A Lower Bound on the Number of Solutions
Under the CSP formulation in Proposition 1, each QC-SC-LDPC design corresponds to an assignment of the $\gamma\kappa$ edge variables $X_{i,j}=(\mathbf{P}(i,j),\mathbf{L}(i,j))$, i.e., a pair $(\mathbf{P},\mathbf{L})$ consisting of an edge-spreading (partition) matrix $\mathbf{P}$ and a lifting-shift matrix $\mathbf{L}$. Therefore, using the quantitative version of the CLLL, we can not only obtain existence results but also establish an explicit lower bound on the number of feasible pairs $(\mathbf{P},\mathbf{L})$ that satisfy all structure-avoidance constraints, thereby quantifying the size of the feasible solution space. The size and structural information of this solution space are crucial for finding codes with good overall performance at lower complexity. In this subsection, we first use the quantitative CLLL to derive a theoretical lower bound on the number of feasible solutions $(\mathbf{P},\mathbf{L})$.
@@
For any QC-SC-LDPC code constructed with respect to a $\gamma\times\kappa$ fully-connected block code base graph, with no parallel edges, under the coupling pattern $\mathbf{a}=(a_{0},a_{1},\ldots,a_{m_{t}})$ and corresponding (non-uniform) distribution $\mathbf{p}=(p_{0},p_{1},\ldots,p_{m_{t}})$, and under the lifting-shift alphabet $\mathbf{b}=(b_{0},b_{1},\ldots,b_{Z-1})=(0,1,\ldots,Z-1)$ with distribution $\mathbf{q}=(q_{0},q_{1},\ldots,q_{Z-1})$. Assume the random construction chooses each $\mathbf{P}(i,j)$ independently with $\Pr[\mathbf{P}(i,j)=\ell]=p_{\ell}$, and chooses each lifting shift $\mathbf{L}(i,j)$ independently with $\Pr[\mathbf{L}(i,j)=b_{t}]=q_{t}$, and $P$ is independent of $L$.
@@
Here, $(\bigcap_{i\in[k]}\bar{H_{i}})$ denotes the number of feasible assignments of $(\mathbf{P},\mathbf{L})$ (equivalently, solutions in the CSP variable set $\chi$) that make $\bigcap_{i\in[k]}\bar{H_{i}}$ true.
@@
See Appendix A of  [10], the notation is consistent with that in  [9]. ∎
@@
${\#\Big(\bigcap_{c_{4}\in\mathcal{C}_{4}}\bar{c_{4}}\Big)}\geq$ $[Z(m_{t}+1)]^{\gamma\kappa}$ (5)
$\times\begin{cases}(1-\frac{2}{\Delta})^{\binom{\gamma}{2}\binom{\kappa}{2}},&{\text{if}}\ I>II,\\[2.0pt] \Big(1-\frac{\gamma\kappa-\gamma-\kappa+1}{4(\gamma\kappa-\gamma-\kappa)}\Big)^{\gamma\kappa},&{\text{otherwise.}}\end{cases}$ (6)
@@
It is straightforward to verify that the expression above takes the form $\exp\big(\Theta(\gamma\kappa)\big)$, implying that the solution space is exponentially large.
@@
In this paper, we employ the concept of equivalence to demonstrate that the decoding threshold and the population of combinatorial objects (such as cycles, absorbing sets [20], and trapping sets [14]) remain invariant under row and column permutations of a binary matrix.
Lemma 2 demonstrates the structural congruence between coupled proto-matrices and their constituent partitioning matrices. Consequently, the design of coupled SC-LDPC codes can be effectively reduced to searching within a compact space of small $(\gamma\times\kappa)$ non-equivalent partitioning matrices.
@@
For any QC-SC-LDPC code constructed with respect to a $\gamma\times\kappa$ fully-connected block code base graph, with no parallel edges, under the coupling pattern $\mathbf{a}=(a_{0},a_{1},...,a_{m_{t}})$ and corresponding probability distribution $\mathbf{p}=(p_{0},p_{1},...,p_{m_{t}})$. If $\{H_{1},...,H_{k}\}$ are a series of avoidable harmful structures in the base graph with dependency graph $G$, and $\{P_{H_{1}},...,P_{H_{k}}\}$ are the probabilities that they are active after the random edge-spreading and lifting process, if
@@
$\#_{N}\left(\bigcap_{i\in[k]}\overline{H_{i}}\right)\geq{}$ $\left(\frac{1}{p_{\max}q_{\max}}\right)^{\gamma\kappa}\left(\frac{1}{\gamma!\kappa!}\right)$ (7)
@@
${\#_{N}(\bigcap_{c_{4}\in\mathcal{C}_{4}}\bar{c_{4}})}\geq\begin{cases}\frac{[Z(m_{t}+1)]^{\gamma\kappa}(1-\frac{2}{\Delta})^{\binom{\gamma}{2}\binom{\kappa}{2}}}{(\gamma!\kappa!)},&{\text{if}}\ I>II\\ {\frac{[Z(m_{t}+1)]^{\gamma\kappa}(1-\frac{\gamma\kappa-\gamma-\kappa+1}{4(\gamma\kappa-\gamma-\kappa)})^{\gamma\kappa}}{(\gamma!\kappa!)},}&{\text{otherwise.}}\end{cases}$
@@ III-B Diversity of Constructive Outputs via MT Entropy
In this subsection, we use an entropy method (via Rényi entropy bounds for the MT output distribution) to obtain a computable lower bound on the number of feasible pairs $(\mathbf{P},\mathbf{L})$ that the algorithm can output. Under the OO-CPO framework, the CPO operation is closely related to the MT algorithm (see Appendix F of [10]).
@@
In the OO-CPO framework, circulant power optimization (CPO) includes the Moser-Tardos (MT) algorithm as a special case.
@@
In the sequel, we show that the MT algorithm can generate a provably large set of distinct feasible solutions. This observation suggests that CPO-based design procedures, such as those in [3, 7, 24], also have significant potential, since Proposition 2 establishes that, within the OO-CPO framework, CPO subsumes the MT algorithm as a special case.
@@
For any QC-SC-LDPC code constructed with respect to a $\gamma\times\kappa$ fully-connected block code base graph,with no parallel edges, under the coupling pattern $\mathbf{a}=(a_{0},a_{1},...,a_{m_{t}})$ and corresponding probability distribution $\mathbf{p}=(p_{0},p_{1},...,p_{m_{t}})$. If $\{H_{1},...,H_{k}\}$ are a series of avoidable harmful structures in the base graph with dependency graph $G$, and $\{P_{H_{1}},...,P_{H_{k}}\}$ are the probabilities that they are active after the random edge-spreading and lifting process, if there is a weighting function $\tilde{\mu}:\mathcal{A}\rightarrow[0,\infty)$ satisfies
@@
for any $i\in[n]$, then the MT Algorithm resamples an event $A\in\mathcal{A}$ at most an expected $\tilde{\mu}(A_{i})$ times before it finds such an evaluation. The number of different edge-spreading and lifting assignments $(\mathbf{P},\mathbf{L})$ that the algorithm can output, denoted as $_{MT}(\bigcap_{i\in[k]}\bar{H_{i}})$, satisfies the following condition:
@@
where $\Delta=(2\gamma-3)(2\kappa-3)$ , $Z$ is the lifting degree. then the number of different edge-spreading and lifting assignments $(\mathbf{P},\mathbf{L})$ that the algorithm can output, denoted as $_{MT}(\bigcap_{c_{4}\in\mathcal{C}_{4}}\bar{c_{4}})$, satisfies the following condition:
@@
Moreover, applying the alternative explicit specialization yields the following lower bound in terms of $l(A)$:
@@
The proof, the comparison of the two lower bounds, and the numerical verification are detailed in Appendices C, D, and E of [10], respectively. ∎
@@
The number $_{MT-N}(\bigcap_{i\in[k]}\bar{H_{i}})$ of non-equivalent edge-spreading and lifting assignments $(\mathbf{P},\mathbf{L})$ satisfying Theorem 3 that can be output by the algorithm is given by:
@@
The number $_{MT-N}(\bigcap_{c_{4}\in\mathcal{C}_{4}}\bar{c_{4}})$ of non-equivalent edge-spreading and lifting assignments satisfying Theorem 3 that can be output by the algorithm is given by:
@@
Let $n:=\gamma\kappa$ and $X:=Z(m_{t}+1)$. Corollary 3 provides an explicit diversity guarantee for the MT resampling procedure: under the stated LLL-type condition, the number of distinct feasible edge-spreading and lifting assignments $(\mathbf{P},\mathbf{L})$ that MT algorithm can output while eliminating all $4$-cycle admits the closed-form lower bounds in (19)–(20), whose dominant term is $X^{n}$ multiplied by a degree-dependent penalty factor. Consequently,
$_{MT}\Bigl(\bigcap_{c_{4}\in\mathcal{C}_{4}}\overline{c_{4}}\Bigr)=X^{n}\exp\!\bigl(-\Theta(n)\bigr)=\exp\!\bigl(\Theta(n)\bigr),$
i.e., the MT output support is exponentially large in the number of edge variables. Moreover, the product-form specialization in (20) is strictly tighter than (19). Corollary 5 further shows that, even after quotienting out row/column-permutation symmetries, the number of non-equivalent MT outputs remains provably large:
$_{MT\text{-}N}\Bigl(\bigcap_{c_{4}\in\mathcal{C}_{4}}\overline{c_{4}}\Bigr)\geq\frac{1}{\gamma!\kappa!}\,\#_{MT}\Bigl(\bigcap_{c_{4}\in\mathcal{C}_{4}}\overline{c_{4}}\Bigr),$
so the algorithm can still generate a large family of genuinely distinct structure-avoiding QC-SC-LDPC designs.